dry air will support a maximum electric field strength of about 2.0x108V/m. Above that value, the field creates enough ionization in the air to make the air a conductor. this allows a discharge or spark that reduces the field. what, then, is the maximum voltage between two parrallel conducting plates separated by 3.5cm of dry air?
1
Expert's answer
2021-04-29T07:44:59-0400

Maximum Field Strength "=2.0\times10^8\space V/m"

Separation between plates, "d=3.5\space cm=3.5\times10^{-2}\space m"

Electric field between two parallel conducting plates,

"E=\dfrac{V}{d}"

"V=E\times d=(2.0\times10^8)\times(3.5\times10^{-2})"

"V=7\times10^{6}\space V"

Maximum Voltage "=7\times10^6\space V"